OpenStudy (anonymous):

Solve this limit approaching 1, algebraically: x^2-1/x^3-1 Please & thank you!

OpenStudy (anonymous):

gimmick is to factor and cancel, then replace x by 1

OpenStudy (anonymous):

Satellite's on it already!

OpenStudy (anonymous):

Can you use L'Hopital on it?

OpenStudy (anonymous):

\[\frac{x^2-1}{x^3-1}=\frac{(x-1)(x+1)}{(x-1)(x^2+x+1)}=\frac{x+1}{x^2+x+1}\]

OpenStudy (anonymous):

replace x by 1 and unless i screwed the pooch on the algebra you get \[\frac{2}{3}\]

OpenStudy (anonymous):

Thank you soooo much, you just saved me from failing my test tomorrow :)

OpenStudy (anonymous):

@mathboy usually these come before derivatives, so long before l'hopital

OpenStudy (anonymous):

@nsl11 if you get \[\frac{0}{0}\] look to factor and cancel

OpenStudy (anonymous):

I got 2/3, thank you!

OpenStudy (anonymous):

and it is not hard to factor if you have polynomials when you are given a 0. factor it out